# Bayesian KalmanNet: Quantifying Uncertainty in Deep Learning Augmented Kalman Filter

### v2

@@ V-A3 Performance Measures
We say that the algorithm is a better uncertainty estimator as its uncertainty prediction (apec along time) is closer to to the true error covariance (eec along time).
@@ V-B Linear Canonical SS Model
When introducing a mismatch in the measurement noise covariance $\mathbf{R}$ and evaluating the filter performance using the anees metric, as shown in Fig. 3(d), it is evident that bkn achieves the most favorable results. bkn provides a credible error covariance estimation, meaning it maintained reliability. The ekf and KalmanNet produce comparable results, indicating that both methods offer similar levels of credibility in their error covariance predictions. In contrast, the rkn struggles to provide credible estimates, leading to unreliable uncertainty predictions.
Turning to the mse results, illustrated in Fig. 3(c), it is clear that the skn outperforms other methods by achieving the lowest estimation error across various measurement noise covariance levels. This indicates that skn consistently provided the most accurate state estimates, regardless of the noise magnitude. The rkn, however, only performed well at higher $r^{2}$ values, suggesting that it struggles in high snr where accurate uncertainty modeling is more critical. All KalmanNet based methods, including bkn, outperform the ekf.
@@
((a)) mse for mismatched $\mathbf{Q}$, all the model aware methods managed to track the state with a very similar mse. the model agnostic rkn has failed to maintaine the same performance
@@ V-C Pendulum SS Model
Next, we consider a pendulum with mass $m=1$ and length $\ell=1$. The pendulum exhibits nonlinear behavior due to the gravitational force $g$ acting on it. We describe the system using the angle $\theta$ and angular velocity $\dot{\theta}$ as the state variables, sampled at interval of $\Delta t$. Let the state vector $\mathbf{x}_{t}$ be defined as:
@@
Here, $\hat{\mathbf{H}}(\mathbf{x}_{t})$ is non-full columns rank. This mean that the frequentist model-based methods are unable to extract an error covariance matrix as explained in Remark 1, ans thus are not tested throughout this scenario.
@@ V-D Navigation SS Model
Figure 6 shows the predicted position x track of each method compared to the ground true trajectory. This figure is a qualitative measure that shows the ability to track the trajectory. As can be seen, the model agnostic rkn failed to track the state along time, whereas all the model aware methods provide a reasonable state estimation.
@@ V-E Inference Time Comparison
The results presented so far evaluate the considered methods in terms of their accuracy and uncertainty estimation performance. As Bayesian deep learning induces excessive latency in inference compared with frequentist models due to its inherent ensembling, we next evaluate inference time. The results in Table II report the inference latency of the different filtering methods across both a linear (canonical) and non-linear (pendulum) scenario. The platform on which all algorithms were tested and runtimes were computed is an Intel(R) Core(TM) i7-10510U CPU @ 1.80GHz 2.30 GHz. In the canonical ss model, the ekf demonstrates the fastest performance of all model aware methods, with the lowest average inference time of $5.279$ milliseconds. The skn, KNet V1, and V2 methods exhibit comparable speeds, with marginal differences around $5.3$ milliseconds. These results suggest that these methods maintain efficient performance close to the traditional ekf, making them suitable for real-time applications where speed is critical.
The rkn achieves an average inference time of just $0.194$ milliseconds. This presumably stems from the efficiency of rnn software implementations in PyTorch. bkn requires substantially more time, with an inference time of $104.88$ milliseconds for the canonical model. This reflects the computational cost associated with more advanced uncertainty quantification and Bayesian learning techniques. While these methods enhance credibility and robustness, they introduce a significant computational burden, making them less ideal for time-sensitive tasks.
@@
Overall, while the bkn offers substantial benefits in terms of credibility and robustness, these come at the cost of longer inference times. For real-time applications, a parallelization could be used to solve this issue, as the different inferences in the bkn are independent and thus could be computed in parallel.
TABLE II: Average Inference Time [msec]

### v3

@@ V-A3 Performance Measures
We say that the algorithm is a better uncertainty estimator as its uncertainty prediction (apec along time) is closer to the true error covariance (eec along time).
@@ V-A4 Mismatch Setup
Our motivation for employing the considered mismatches, imposed on all algorithms that require knowledge of the ss model, is to evaluate sensitivity to incorrect modeling assumptions. Specifically, the noise mismatches (which affect just the ekf) serve to highlight the behavior of uncertainty quantification under severe mismatches, while the last model mismatch represents a realistic setup in which a target follows a more complicated state evolution model compared to the one used by the tracking algorithm.
@@ V-B Linear Canonical SS Model
When introducing a mismatch in the measurement noise covariance $\mathbf{R}$ and evaluating the filter performance using the anees metric, as shown in Fig. 3(d), it is evident that bkn achieves the most favorable results. bkn provides a credible error covariance estimation, meaning it maintained reliability. The ekf and KalmanNet produce comparable results, indicating that both methods offer similar levels of credibility in their error covariance predictions. In contrast, the rkn struggles to provide credible estimates, leading to unreliable uncertainty predictions. Comparing Figs. 3(b) and 3(d) shows that frequentist methods can also achieve strong performance, particularly in scenarios where uncertainty can be reliably estimated. However, in cases where uncertainty estimation is inherently more challenging, e.g., where inaccuracies directly impact the observed data—frequentist methods relying on internal features may struggle. In such settings, Bayesian KalmanNet provides more robust and calibrated uncertainty estimates by explicitly modeling parameter uncertainty through ensemble predictions. Overall, both variants exhibit stable and credible uncertainty estimates across different SNR levels, and the observed differences are relatively minor.
Turning to the mse results, illustrated in Fig. 3(c), it is clear that the skn, which is particularly designed for settings with imbalances between state evolution and measurement noises [23], outperforms other methods by achieving the lowest estimation error across various measurement noise covariance levels. This indicates that skn consistently provided the most accurate state estimates, regardless of the noise magnitude. The rkn, however, only performed well at higher $r^{2}$ values, suggesting that it struggles in high snr with limited data sets, where accurate uncertainty modeling is more critical. All KalmanNet based methods, including bkn, outperform the ekf. Comparing the mse results in Fig. 3(c) and the uncertainty results in Fig. 3(d) indicates on the trade-off emerging by the need to balance both objectives during training. In such multi-objective settings, improving one aspect-—e.g., achieving lower MSE-—may slightly compromise the precision of uncertainty estimates, sometimes somewhat compromise the precision of uncertainty estimates, and in some cases hardly affect the latter.
@@
((a)) mse for mismatched $\mathbf{Q}$, all the model aware methods managed to track the state with a very similar mse. The model agnostic rkn has failed to maintain the same performance.
@@ V-C Pendulum SS Model
Next, we consider a pendulum with mass $m=1$ and length $\ell=1$. The pendulum exhibits nonlinear behavior due to the gravitational force $g$ acting on it. We describe the system using the angle $\theta$ and angular velocity $\dot{\theta}$ as the state variables, sampled at interval of $\Delta t=0.01$. Let the state vector $\mathbf{x}_{t}$ be defined as:
@@
Here, $\hat{\mathbf{H}}(\mathbf{x}_{t})$ is non-full columns rank. This means that the frequentist model-based methods are unable to extract an error covariance matrix as explained in Remark 1, and thus are not tested throughout this scenario.
@@
TABLE II: Performance metrics at snr $101/r^{2}=$ 10 [dB] with mismatched $\mathbf{R}$, Pendulum non-linear ss model.
Metric EKF SKN RKN BKN
MSE [dB] 23.8933 -7.5536 -7.7836 -9.2741
log(ANEES) 4.1182 1.2452 0.2234 -0.0834
@@ V-D Navigation SS Model
We generate a set of 100 trajectories, with 5 used for validation and 10 reserved for testing. The sampling interval parameter $\Delta t$ is normalized to unity.
@@
Fig. 6 shows the predicted position x track of each method compared to the ground truth trajectory. This figure is a qualitative measure that shows the ability to track the trajectory. As can be seen, the model agnostic rkn failed to track the state along time, whereas all the model aware methods provide a reasonable state estimation.
@@ V-E Inference Time Comparison
The results presented so far evaluate the considered methods in terms of their accuracy and uncertainty estimation performance. As Bayesian deep learning induces excessive latency in inference compared with frequentist models due to its inherent ensembling, we next evaluate inference time. The results in Table III report the inference latency of the different filtering methods across both a linear (canonical) and non-linear (pendulum) scenario. The platform on which all algorithms were tested and runtimes were computed is an Intel(R) Core(TM) i7-10510U CPU @ 1.80GHz 2.30 GHz. For bkn (bkn), The latency values reflect the total execution time required to process all $J$ realizations sequentially, i.e., the actual inference latency experienced in a serial execution setting. In the canonical ss model, the ekf demonstrates the fastest performance of all model aware methods, with the lowest average inference time of $5.279$ milliseconds. The skn, KNet V1, and V2 methods exhibit comparable speeds, with marginal differences around $5.3$ milliseconds. These results suggest that these methods maintain efficient performance close to the traditional ekf, making them suitable for real-time applications where speed is critical.
The rkn achieves an average inference time of just $0.194$ milliseconds. This presumably stems from the efficiency of rnn software implementations in PyTorch. bkn requires substantially more time, with an inference time of $104.88$ milliseconds for the canonical model. This reflects the computational cost associated with more advanced uncertainty quantification and Bayesian learning techniques. While these methods enhance credibility and robustness, they introduce a significant computational burden, making them less suited for time-sensitive tasks.
@@
Overall, while the bkn offers substantial benefits in terms of credibility and robustness, these come at the cost of longer inference times. For real-time applications, a parallelization could be used to solve this issue, as the different inferences in the bkn are independent and thus could be computed in parallel. While there is no consistent algorithm that achieves the best performance in both accuracy and uncertainty throughout the various settings considered in our numerical study, it is clearly demonstrated that incorporating domain knowledge notably facilitates uncertainty extraction. Moreover, it is systematically shown that bkn enables dnn-aided kf to both extract the state and predict the error accurately in nonlinear, possibly mismatched ss models, and that its reliability is maintained across all the considered settings.
TABLE III: Average Inference Time [msec].